OpenStudy (anonymous):

What is the specific heat of an unknown substance if 2000 J of energy are required to raise the temperature of 4 grams of the substance 5 degrees Celsius? Help!!!

OpenStudy (anonymous):

Q = mc(dT) c = Q/m(dT) Q = 2000 J m = 0.004 kg dT = 5 c = ? c = (2000)/[(0.004)(5)] c = 2000 / 0.02 Therefore, specific heat capacity c = 100,000 J/kg = 100 kJ/kg